Solve the initial value Problem
dN/dt = τ (N − N^2/q) N(0) = No, where τ, q are positive. Hence or otherwise, find the solution of the Verhulst model
when t → ∞
1
Expert's answer
2020-08-05T17:55:09-0400

Given equation is

"\frac{dN}{dt} = \tau (N - \frac{N^2}{q})"

Integrating the above equation

"\frac{dN}{(\frac{N^2}{q} - N)} = - \tau dt"


Integrating it,

"\int \frac{dN}{(\frac{N^2}{q} - N)} = - \int\tau dt"

Solving it we get,

"ln|1-\frac{q}{N}| = -\tau t + C"


Putting boundary equations, N(0) = N0

Then we get

"ln|1-\frac{q}{N_0}| = C"



Hence, we can say that

"ln|1-\frac{q}{N}| = -\tau t + ln|1-\frac{q}{N_0}|"


It can be written as

"ln|\frac{N(N_0-q)}{N_0(N-q) }| = \tau t"